A projectile is fired at 27.0 m/s in a direction 65⁰ above the horizonal from a rooftop of height 40.0 m. Use energy considerations to find: (a) the speed with which it lands on the ground; (b) the height at which its speed is 15.0 m/s.
Answer #1

a)

By Conservation of energy

(1/2)mVi2+mgh =(1/2)mVf2

(1/2)*272 + 9.81*40 = (1/2)*Vf2

Vf = 38.908 m/s

b)

By Conservation of energy

(1/2)mVi2+mgh =(1/2)mVf2+mghf

(1/2)*272 + 9.81*40 = (1/2)*152+9.81*hf

hf = 65.688 m